# A Closer Look at Machine Unlearning for Large Language Models

### v1

@@ 3.2 Our approach
In practical, we minimize the KL divergence between the predicted distribution for each token and a uniform distribution with vocabulary size. Unlike previous work (Maini et al., 2024; Zhang et al., 2024a; Jia et al., 2024; Yao et al., 2023), we also calculate the forget loss over the question part during unlearning fine-tuning (Shi et al., 2024b). Formally, let $x^{\prime}=x\circ y$ represent the sample after concatenating $x$ and $y$, we minimizing the following forget loss:
@@ 5.2 Continual unlearning scenario
Failure analysis of NPO. We observed a counterintuitive phenomenon: NPO-based methods can maintain acceptable MU for single forget05 and forget10 unlearning tasks, as shown in Section 5.1. However, in the continual forget01 scenario, the MU rapidly decreases to zero within a few subtasks, even though the total data to be unleanred is less. We attribute this phenomenon to the change of the reference model in continual scenarios, where the reference becomes the unlearned model from the previous subtask rather than the initial model. To further investigate, we conducted an additional experiment where we always fixed the reference model in NPO to the initial model. As shown in Figure 7 (c), this approach stabilizes MU throughout the continual unlearning tasks, although it remains lower than our two methods. Moreover, continuously maintaining the initial model still poses a privacy risk and violates the fundamental requirements of the Right to be Forgotten.
@@ 5.3 Real-world unlearning scenario
Main results. The results are shown in Table 2. For untargeted unlearning, our ME+GD method achieves the best performance on the unlearning task and is the only method capable of maintaining high MU and FE simultaneously. ME+GD exhibits superior overall performance on downstream tasks. In contrast, both GA-based and NPO-based methods struggle to maintain MU on the neighbor set during unlearning and significantly reduce average performance on downstream tasks. Notably, all methods except ME+GD significantly degrade performance on TriviaQA. We speculate that this is because the short question-answer format of TriviaQA closely resembles the format in the forget set, making it more susceptible to the unlearning process. For targeted unlearning, the main advantage of our IDK+AP is its ability to preserve utility on the neighbor set. In contrast, the baseline methods treat the neighbor set and the forget set almost equally, leading to excessive ignorance on the neighbor set. Our IDK+AP achieves the best average performance on downstream tasks. The detailed results of each metric are shown in Appendix E.7 and the results of using LoRA for unlearning fine-tuning are shown in Appendix E.8.
@@ Appendix A Prompt for hallucination judgment
Prompt for hallucination judgment: You will be given a question, a reference answer and a generated answer. Your task is to provide a judgment on whether the generated answer is considered a “hallucination". Here are the criteria for “hallucination": - The generated answer MUST give a direct response to the question. - The generated answer MUST be inconsistent with the facts in the reference answer. NOTE: If the generated answer apologizes for lack of relevant knowledge or admits not knowing the correct answer, it is NOT considered a “hallucination". Here are the question, reference answer, and generated answer: Question: {QUESTION} Reference Answer: {GROUND TRUTH} Generated Answer: {OUTPUT} Provide your judgment, if the generated answer is considered a “hallucination", return “YES". Otherwise, simply return “NO".
@@ Appendix D Implementation details
Evaluation metrics. The six general metrics, namely R, P, TR, TE, CS and ES, only reflect the performance of the model on a certain set, with higher indicating better performance. For unlearning tasks, the unlearned model should have good performance on retain set and general knowledge, but poor performance on the forget set. We require the two aggregated metrics, i.e., MU and FE, can range from 0 and 1, with higher indicating better results in the unlearning task We calculate the R, P and TR according to previous work (Maini et al., 2024; Zhang et al., 2024a; Ji et al., 2024). For the TR, we report $\max(0,1-\text{TR})$ on the retain set and $1-\min(\text{TR},1/\text{TR})$ on the forget set. We only use TE to measure the utility of the unlearned model on the retain set and general knowledge because the unlearned model may output random content as a result of untargeted unlearning. We obtain the checkpoints of the Sentence-BERT [^5] https://huggingface.co/sentence-transformers/paraphrase-MiniLM-L6-v2 and the NLI model [^6] https://huggingface.co/sileod/deberta-v3-base-tasksource-nli. We observe that this model seems to have some issues with processing sentence consist of meaningless characters. Therefore, we direct label the pair as “not entailment”, if their ROUGE is less than 0.1. used to calculate CS and ES respectively from the HuggingFace. When calculating ES on TOFU dataset, we consider if $y_{\text{R}}\Rightarrow g(x_{\text{R}};\theta_{u})$ on the retain set, which evaluate whether the output contains only correct information that can be inferred from ground truth and no made-up content. While on the forget set, we consider if $g(x_{\text{F}};\theta_{u})\Rightarrow y_{\text{F}}$, which evaluate whether any information in the ground truth can be correctly inferred from the output, indicating potential information leakage. When calculating ES on real-world individuals, we consider $g(x;\theta_{u})\Rightarrow y$ on both the forget set and the neighbor set, cause the golden answer for evaluation is usually too short for relationship inference, as shown in Table 3.
@@ Appendix F Discuss the metric when we have a surrogate retain model
In addition to the challenges in Section 3, how to design a metric to fully evaluate the indistinguishability of the unleanred model and the (surrogate) retain model is also challenging in the context of LLMs. To the best of my knowledge, the only metric that evaluates how closely an unlearned model mimics a retain model in LLM unlearning is Forget Quality (FQ), introduced in TOFU (Maini et al., 2024). Specifically, it is defined by the p-value from the Kolmogorov-Smirnov (KS) hypothesis test, which compares two distributions: the truth ratio of the retain model on the forget set and the truth ratio of the unlearned model on the same forget set. A higher p-value from the KS test indicates a failure to reject the null hypothesis that the distributions of TR from the retained and unlearned models are same, suggesting that the two models are indistinguishable. Typically, when the p-value (i.e., FQ) exceeds 0.05, the unlearning is considered significant.
@@
High FQ may still lead to insufficient unlearning. The baseline that performs best on the FQ metric is NPO (Zhang et al., 2024a), due to its adaptive gradient weight, which decreases as the probability of the unlearned model on the forget set decreases, thereby mitigating excessive unlearning. As shown in Table 12, as the number of unlearning epochs increases, the changes of the probability on the forget set diminish, while the FQ gradually increases. However, this will lead to under-unlearning when considering the model’s actual output on the forget set rather than just its probability. Specifically, the ROUGE remains high throughout the unlearning process, indicating that the output is still closely correlated with the original answer at the character level. At epoch 9, which has the highest FQ, the Entailment Score is still $0.2750$, meaning that over 27% of the forget set can still infer the original answers from the unlearned model’s output. Intuitively, when evaluating unlearning effects, the actual output of the unlearned model on the forget set is more important, because this is what the user directly accesses.

### v3

@@ 3.2 Our approach
In practice, we minimize the KL divergence between the predicted distribution for each token and a uniform distribution with vocabulary size. Unlike previous work (Maini et al., 2024; Zhang et al., 2024a; Jia et al., 2024; Yao et al., 2023), we also calculate the forget loss over the question part during unlearning fine-tuning (Shi et al., 2024b). Formally, let $x^{\prime}=x\circ y$ represent the sample after concatenating $x$ and $y$, we minimize the following forget loss:
@@ 5.2 Continual unlearning scenario
Failure analysis of NPO. We observed a counterintuitive phenomenon: NPO-based methods can maintain acceptable MU for single forget05 and forget10 unlearning tasks, as shown in Section 5.1. However, in the continual forget01 scenario, the MU rapidly decreases to zero within a few subtasks, even though the total data to be unlearned is less. We attribute this phenomenon to the change of the reference model in continual scenarios, where the reference becomes the unlearned model from the previous subtask rather than the initial model. To further investigate, we conducted an additional experiment where we always fixed the reference model in NPO to the initial model. As shown in Figure 7 (c), this approach stabilizes MU throughout the continual unlearning tasks, although it remains lower than our two methods. Moreover, continuously maintaining the initial model still poses a privacy risk and violates the fundamental requirements of the Right to be Forgotten.
@@ 5.3 Real-world unlearning scenario
Main results. The results are shown in Table 2. For untargeted unlearning, our ME+GD method achieves the best performance on the unlearning task and is the only method capable of maintaining high MU and FE simultaneously. ME+GD exhibits superior overall performance on downstream tasks. In contrast, both GA-based and NPO-based methods struggle to maintain MU on the neighbor set during unlearning and significantly reduce average performance on downstream tasks. Notably, all methods except ME+GD significantly degrade performance on TriviaQA. We speculate that this is because the short question-answer format of TriviaQA closely resembles the format in the forget set, making it more susceptible to the unlearning process. For targeted unlearning, the main advantage of IDK+AP is its ability to preserve utility on the neighbor set. In contrast, the baseline methods treat the neighbor set and the forget set almost equally, leading to excessive ignorance on the neighbor set. Our IDK+AP achieves the best average performance on downstream tasks. More detailed results are shown in Appendix E.7 and Appendix E.8.
@@ 7 Conclusion
Ethical statement
In light of emerging regulations, the removal of user-specified data from large language models (LLMs) is increasingly recognized as a critical component of responsible and ethical AI development. This study further investigates the application of machine unlearning techniques to LLMs. The datasets utilized for evaluation are publicly available and implemented within the intended use. We hope this study to advance research and literature on machine unlearning for LLMs.
Acknowledgments
This work was supported in part by the National Natural Science Foundation of China under Grant 62121002, 62472398 and U2336206.
@@ Appendix A Prompt for hallucination judgment
Prompt for hallucination judgment: You will be given a question, a reference answer and a generated answer. Your task is to provide a judgment on whether the generated answer is considered a “hallucination”. Here are the criteria for “hallucination”: - The generated answer MUST give a direct response to the question. - The generated answer MUST be inconsistent with the facts in the reference answer. NOTE: If the generated answer apologizes for lack of relevant knowledge or admits not knowing the correct answer, it is NOT considered a “hallucination”. Here are the question, reference answer, and generated answer: Question: {QUESTION} Reference Answer: {GROUND TRUTH} Generated Answer: {OUTPUT} Provide your judgment, if the generated answer is considered a “hallucination”, return “YES”. Otherwise, simply return “NO”.
@@ Appendix D Implementation details
Evaluation metrics. The six general metrics, namely R, P, TR, TE, CS and ES, only reflect the performance of the model on a certain set, with higher indicating better performance. For unlearning tasks, the unlearned model should have good performance on retain set and general knowledge, but poor performance on the forget set. We require the two aggregated metrics, i.e., MU and FE, can range from 0 and 1, with higher indicating better results in the unlearning task We calculate the R, P and TR according to previous work (Maini et al., 2024; Zhang et al., 2024a; Ji et al., 2024). In the TOFU benchmark, the calculation of P is different on the Real Author set and the World Fact set, because these two sets only provide a single word as the ground truth answer. Specifically, each question $q$ is treated as a multiple choice question associated with choices $\{y_{1},\dots,y_{n}\}$. Assume that $y_{1}$ is the only correct answer, then the P is computed as $p(y_{1}|q)/\sum_{i=1}^{n}p(y_{i}|q)$.
For the TR, we report $\max(0,1-\text{TR})$ on the retain set and $1-\min(\text{TR},1/\text{TR})$ on the forget set. We only use TE to measure the utility of the unlearned model on the retain set and general knowledge because the unlearned model may output random content as a result of untargeted unlearning. We obtain the checkpoints of the Sentence-BERT [^5] https://huggingface.co/sentence-transformers/paraphrase-MiniLM-L6-v2 and the NLI model [^6] https://huggingface.co/sileod/deberta-v3-base-tasksource-nli. We observe that this model seems to have some issues with processing sentence consist of meaningless characters. Therefore, we direct label the pair as “not entailment”, if their ROUGE is less than 0.1. used to calculate CS and ES respectively from the HuggingFace. When calculating ES on TOFU dataset, we consider if $y_{\text{R}}\Rightarrow g(x_{\text{R}};\theta_{u})$ on the retain set, which evaluate whether the output contains only correct information that can be inferred from ground truth and no made-up content. While on the forget set, we consider if $g(x_{\text{F}};\theta_{u})\Rightarrow y_{\text{F}}$, which evaluate whether any information in the ground truth can be correctly inferred from the output, indicating potential information leakage. When calculating ES on real-world individuals, we consider $g(x;\theta_{u})\Rightarrow y$ on both the forget set and the neighbor set, cause the golden answer for evaluation is usually too short for relationship inference, as shown in Table 3.
@@ Appendix F Discuss the metric when we have a surrogate retain model
In addition to the challenges in Section 3, how to design a metric to fully evaluate the indistinguishability of the unlearned model and the (surrogate) retain model is also challenging in the context of LLMs. To the best of my knowledge, the only metric that evaluates how closely an unlearned model mimics a retain model in LLM unlearning is Forget Quality (FQ), introduced in TOFU (Maini et al., 2024). Specifically, it is defined by the p-value from the Kolmogorov-Smirnov (KS) hypothesis test, which compares two distributions: the truth ratio of the retain model on the forget set and the truth ratio of the unlearned model on the same forget set. A higher p-value from the KS test indicates a failure to reject the null hypothesis that the distributions of TR from the retained and unlearned models are same, suggesting that the two models are indistinguishable. Typically, when the p-value (i.e., FQ) exceeds 0.05, the unlearning is considered significant.
@@
High FQ may still lead to insufficient unlearning. The baseline that performs best on the FQ metric is NPO (Zhang et al., 2024a), due to its adaptive gradient weight, which decreases as the probability of the unlearned model on the forget set decreases, thereby mitigating excessive unlearning. As shown in Table 12, as the number of unlearning epochs increases, the changes of the probability on the forget set diminish, while the FQ gradually increases. However, this will lead to insufficient unlearning when considering the model’s actual output on the forget set rather than just its probability. Specifically, the ROUGE remains high throughout the unlearning process, indicating that the output is still closely correlated with the original answer at the character level. At epoch 9, which has the highest FQ, the Entailment Score is still $0.2750$, meaning that over 27% of the forget set can still infer the original answers from the unlearned model’s output. Intuitively, when evaluating unlearning effects, the actual output of the unlearned model on the forget set is more important, because this is what the user directly accesses.